# Exploring Prompting Methods for Mitigating Class Imbalance through Synthetic Data Generation with Large Language Models

### v2

@@ 2.5 Overview of tabular data generation process
Synthetic data generation starts with creating a prompt template (see Appendix I.5 for examples). This template determines the target class and the number of sets, groups, and samples per group to construct demonstrations. If numerous categorical variables share the same unique values, we apply unique variable mapping to the dataset. Next, we randomly sample data examples from the dataset to fill in this template. Based on the provided examples, the LLM generates new data, which is then converted back into a tabular format. Instances with categorical variable values not present in the training data are discarded to maintain data authenticity. This process is repeated until the desired number of samples is obtained. Each iteration involves newly sampled examples from the original dataset, allowing the LLM to encounter various combinations and produce diverse samples. Although each individual generation is based on the distribution of the given subset, cumulatively, the iterations enable the model to explore the entire range of the original data. This method ensures that the output generated by the LLM comprehensively reflects the original data characteristics (See Appendix G).
Table 1: Comparison of machine learning classification performance when synthetic data are added to the original data. The results are averaged across four classifiers, with each model run five times. Complete results, including all baselines and standard deviation values, are provided in Appendix K.
@@
3 Experiments
@@
Datasets    We use six real-world classification datasets from diverse domains to evaluate our method: Travel, Sick, HELOC, Income, Diabetes, and Thyroid. The Thyroid dataset is released after the training cut-off date for GPT-3.5-turbo-0613, which provides a more rigorous validation of our approach on the completely unseen dataset. For binary classification datasets, the minor class is set to be one. All duplicate data samples are removed from the datasets. Subsequently, we randomly split them into 80% training and 20% test sets. More details for the datasets are provided in Appendix I.2.
Evaluation measure    We evaluate our generation methodology from three critical aspects: ML classification performance, data generation success rate and stability with LLM in-context learning, and validity of generated data in terms of feature correlation and value range. For the classification task, we measure the F1 score, sensitivity, specificity, and balanced accuracy (BAL ACC). To measure the feature correlation, we use Pearson correlation and Cramér’s V correlation.
@@
Baseline models    In our study, we compare our method with various generative models for tabular data generation, including TVAE [30], CopulaGAN [21], CTGAN [30], CTAB-GAN [33], CTAB-GAN+ [34], GReaT [4], and TabDDPM [18]. This comprehensive evaluation assesses the effectiveness of our approach across different data synthesis paradigms. To compare the prompting method, we also use the prompts used in CuratedLLM [25] and LITO [31].
Experimental details    Our method utilizes the OpenAI GPT-3.5 models, the GPT-3.5-turbo-0613 and GPT-3.5-turbo-16k-0613 versions, Mistral-7b-v0.1 [14], and Llama-2-7b [27]. Unless stated otherwise, four machine learning classification models are used: XGBoost [8], CatBoost [23], LightGBM [16], and Gradient boosting classifier [11]. We execute five independent runs for each model and report the average values. Additional details are provided in Appendix I.
@@
Table 2: Comparison of classification performance using synthetic data generated by open-source LLMs, Llama2 and Mistral, with the proposed prompting strategy, utilizing the Gradient Boosting classifier. We report the average performances for five runs. #syn denotes the number of synthetic samples integrated into the original dataset.
@@ 3.1 Machine learning classification performance using the synthetic data
We assess the quality of the synthetic data samples by comparing the machine learning classification performance averaged over 20 experiments. As shown in Table 1, our method achieves state-of-the-art F1 scores and balanced accuracy across all six datasets, surpassing the baselines that require model training. Notably, our method is the only one that consistently outperforms the original data in both F1 score and balanced accuracy across all six datasets. Other methods often result in worse performance than the original data, indicating that the data generated by baseline methods disrupts the original data distribution. These findings underscore the robustness and effectiveness of our method, affirming its superiority over baseline models, even for challenging datasets.
Moreover, baseline methods such as GReaT and TabDDPM exhibit significantly lower sensitivity compared to specificity, particularly for Travel and Income. Their high balanced accuracy is due to high specificity, but their severely low sensitivity indicates a failure to generate appropriate minority class data. In contrast, our method significantly improves sensitivity and achieves a balance between sensitivity and specificity while also attaining the best-balanced accuracy and F1 score compared to the baselines. For example, in Travel, our method achieves a sensitivity of 78%, which is 19%p higher than the second-best method, GReaT, at 58.8%. In Income, our method shows 66.45% sensitivity, surpassing the second-best score of 60.69%. These results highlight the effectiveness of our method in accurately representing minority classes in tabular datasets.
Overall, our findings demonstrate that our approach is generally applicable across various imbalanced tabular datasets and excels in generating high-quality samples that improve ML classification performance. For detailed confusion matrix results, see Appendix J, and for further analyses exploring the impact of adding synthetic data only to the minor class or using synthetic data alone, see Appendix A.
@@ 3.2 Open-source LLMs
Table 4: Comparison of token efficiency and generation stability on the Income dataset.
@@
Table 5: Comparison of classification performance on task specification elements in prompt design, averaged across four classifiers: XGBoost, CatBoost, LightGBM, and Gradient Boosting classifier. Methods marked with an asterisk (*) use prompts as proposed in the respective papers.
@@
Figure 4: The difference between Cramér’s V correlation matrices of real and synthetic datasets for categorical variables in the Travel dataset. More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue. Black indicates where the correlation is not measured since only one unique value is generated for those variables.
@@ 3.3 Exploring optimal prompt design through ablation studies
We investigate key prompt design elements focusing on ML performance and generation efficiency, as detailed in Tables 3 and Table 4. Token efficiency is measured by the number of input tokens required for a fixed number of input samples. LLM generation efficiency is evaluated based on the average number of samples generated per iteration and the success rate of generating at least one valid data sample per inference. This evaluation is based on 100 iterations, each with 20 random input samples.
@@
Our results indicate that CSV-style prompting generally outperforms sentence-style prompting in ML performance with the same number of input samples. While generating data for one class at a time yields good results for the Sick and Income datasets, it results in a low F1 score of 58.67% on the Travel dataset. However, generating data for multiple classes simultaneously with balanced samples improves this score to 70.37%. Furthermore, applying unique variable mapping consistently enhances F1 scores on the Sick dataset. The combined approach of random mapping and grouping achieves higher F1 score of 91.95% compared to 86.60% with random mapping but without grouping, underscoring the positive impact of this combined approach. These findings indicate that prompt design significantly influences the quality of generated data, emphasizing the need for carefully tailored prompts to generate high-quality synthetic tabular data.
@@
For the Travel and Income datasets, performance without grouping is slightly better than with grouping, but the differences are within 0.5%p, making their performance on par. However, these show significant differences in LLM generation efficiency. Without grouping, an average of 0.88 samples are generated per iteration, whereas grouping increases this to an average of 8.6 samples per iteration. Repeating the set twice within the prompt further improves generation efficiency, boosting the average output to over 10 samples and the success rate to 94%, by creating input patterns that LLMs can mimic to generate data. Unique variable mapping also significantly enhances LLM generation efficiency and success rates, as detailed in Appendix B.
@@
Furthermore, we conduct an ablation study on the task specification elements, as shown in Table 5. Our approach of triggering completion significantly enhances overall ML performance compared to the instruction-based prompting methods. Interestingly, adding class distinctions within the prompt achieves the best sensitivity, indicating that the distinction facilitates better generation for minority classes. Providing variable descriptions improves balanced accuracy but reduces the F1 score, suggesting that this element may be optional rather than a definitive improvement. For a comparison of generation efficiency using these methods, see Appendix B.
3.4 Analysis on feature correlation
@@
We compare the feature correlations of the categorical variables between original and synthetic data, as illustrated in Fig. 4. Our method, utilizing Mistral and Llama2, produces data that most closely mirrors the original feature correlations for both minor and major classes. Moreover, when compared to the ablated versions, our grouping method significantly contributes to preserving feature correlation for the minority class. For additional feature correlation results on the Sick dataset, see Appendix C.
@@ 4 Conclusion
This study demonstrates the effectiveness of using LLMs for tabular data generation and introduces an optimal solution for prompt design to generate realistic and reliable data without additional training, particularly addressing class imbalance. Our extensive evaluations demonstrate the robustness and effectiveness of our method, which significantly improves ML classification performance on the six real-world tabular datasets. This highlights its practicality and potential impact on improving model performance in various application scenarios, contributing significantly to the field.
@@
  • Appendix A presents more experimental results on the machine learning classification for imbalanced classes.
  • Appendix B offers complete additional ablation study results.
  • Appendix C provides feature correlation results for the Sick dataset.
  • Appendix D studies the impact of sample size on the machine learning performance.
  • Appendix E comprehensively analyzes the experiments conducted on the toy dataset.
  • Appendix F discusses about the broader impacts of our research.
  • Appendix G outlines limitations of our work.
  • Section H provides related work.
  • Appendix I provides comprehensive information on the datasets, comparison baselines, implementation details, and prompt design and examples.
  • Appendix J presents confusion matrix results, highlighting the effectiveness of our method in addressing minor class problems.
  • Appendix K shows the complete results of Table 1.
@@ Appendix A Additional experimental results
Our experiments with synthetic tabular data for imbalanced classes follow three main approaches: (1) Adding synthetic data to both minor and major classes of the existing dataset and evaluating machine learning classification performance. The results of this approach are presented in Table 1 of the main manuscript. (2) Adding synthetic data only to the minor class, with detailed results provided in Appendix A.1. (3) Evaluating classification performance using only synthetic data without the original dataset, as discussed in Appendix A.2.
Table 6: Comparison of classification performance for imbalanced data by generating synthetic data for the minority class to match the size of the majority class. The average performances of the Gradient Boosting Classifier over five runs are reported.
@@
A.1 Comparative analysis with adding the new dataset only to the minor class
One of the major advantages of our method is its ability to generate high-quality synthetic data for minor classes, even with limited samples. Similar to oversampling techniques like SMOTE, we add synthetic dataset only to the minor class and evaluate machine learning classification performance. To balance the ratio of major to minor classes, we add only minor class samples to the original data, then measure the classification performance of the machine learning model. For this task, we additionally investigate the effectiveness of our method compared to SMOTE and SMOTENC [6]. As shown in Table 6, adding minor class samples generated from our method leads to higher F1 scores and balanced accuracy on both Travel and Income datasets, compared to when samples generated by other baselines are added. In the Travel dataset, where the number of minor class samples is substantially low, baseline models, especially, CopulaGAN and CTGAN, demonstrate markedly low sensitivity, indicating their failure to accurately learn the minor class data distribution. In contrast, our method demonstrates its efficacy in producing high-quality data for minor classes on the Travel dataset. Compared to the Original, it significantly improves the F1 score by 12.13%p, underscoring its superior performance of our approach.
A.2 Comparative analysis of classification performance using only synthetic data
Table 7: Comparison of classification performance using only synthetic data for training. We report average values for five runs employing Gradient Boosting Classifier.

### v4

@@ 2.5 Overview of tabular data generation process
The synthetic data generation process begins by creating a prompt template and determining whether the unique variable mapping is needed. The prompt template specifies parameters such as the target class, sample size per group, the number of groups, and the number of sets to construct demonstrations. We then randomly sample data examples from the dataset to populate the template, ensuring that examples do not overlap within each prompt to maintain diversity. Using the provided examples, the LLM generates new data in a CSV format, which is subsequently converted into a structured tabular format. Instances containing categorical variable values not present in the original dataset are discarded to maintain authenticity. This generation process is repeated until the desired sample size is achieved, with new examples sampled with replacement in each iteration. This approach enables the LLM to encounter diverse combinations of real samples and produce rich synthetic samples (see Appendix B.1 for analysis). Although each individual generation reflects the distribution of the provided subset, successive iterations enable the LLM to cover a broader range of the original data cumulatively. This iterative method ensures that the output generated by the LLM comprehensively reflects the characteristics of the original data. Final prompting examples are available in Appendix C.5
3 Experiments
This section describes the experimental setup and provides a comprehensive evaluation of our method, assessing its effectiveness across diverse real-world datasets.
Table 1: Comparison of ML classification performance when synthetic data are added to the original dataset. Results are averaged across four classifiers, with each model run five times. Complete results, including all baselines and standard deviation values, are provided in Table 19.
@@ 3 Experiments
Summary    We conduct three unique experiments to analyze the utility of our method in enhancing ML classification performance: augmenting the original dataset with generated data (Tables 119, Figs. 214), augmenting only the minority class similar to SMOTE (Table 6), and using only generated data (Table 7). Additionally, we evaluate our approach using three LLMs: Mistral, Llama2, and gpt3.5-turbo (Table 2, Figs. 45). We perform ablations on prompt elements to compare classification performance (Tables 35) and conduct unique analyses of token usage and LLM generation efficiency (Tables 489). To examine feature correlations, we separately analyze minority and majority classes, comparing results across all prompt variations (Figs. 45). We further investigate how varying the number of generated samples affects classification performance (Figs. 67). Using a toy dataset, we analyze how different prompts influence the accuracy of generated data distributions (Figs. 19). Lastly, we explore the sampling of input examples and their corresponding outputs in LLMs using the toy dataset, providing insights into the variability and reliability of generated data (Figs. 1011).
3.1 Experimental setup
@@
Datasets    We evaluate our method using six real-world public tabular classification datasets from diverse domains: Travel (Marketing), Sick (Healthcare), HELOC (Finance), Income (Social science), Diabetes (Healthcare), and Thyroid (Healthcare). The Thyroid dataset, released after the training cut-off date for GPT-3.5-turbo-0613, provides a more rigorous validation of our approach on completely unseen data. For binary classification datasets, the minority class is designated as one. All duplicate data samples are removed from the datasets. Each dataset is randomly split into 80% training and 20% test sets. We retain the original data, including missing or noisy features. The exception is the Sick dataset, where we follow the source’s method. Further details are provided in Appendix C.2.
Evaluation measure    We evaluate our method based on ML classification performance, generation efficiency, feature correlation, and analysis of generated data distribution. For classification, we report F1 score, sensitivity, specificity, and balanced accuracy (BAL ACC). Feature correlation is measured using Pearson correlation for numerical variables and Cramér’s V correlation for categorical variables.
@@
Baseline models    In our study, we compare our method with various generative models for tabular data, including SMOTE [5], SMOTENC [5], TVAE [35], CopulaGAN [22], CTGAN [35], CTAB-GAN [39], CTAB-GAN+ [40], GReaT [4], and TabDDPM [19]. To compare prompting methods, we also use the prompts from CuratedLLM [27] and LITO [37].
Experimental details    Our method utilizes the GPT-3.5 models (GPT-3.5-turbo-0613 and GPT-3.5-turbo-16k-0613), Mistral-7b-v0.1 [15], and Llama-2-7b [29]. Unless stated otherwise, we use the GPT-3.5 model for our method. The number of synthetic data samples is based on the size of the original datasets. Across all experiments, unless stated otherwise, we report results from four top-performing ML classifiers: XGBoost [7], CatBoost [24], LightGBM [17], and Gradient boosting classifier [10], known for their strong performance, often surpassing recent deep learning models on tabular datasets. Each classifier is executed in five independent runs, with results averaged over a total of 20 runs to ensure robustness. Further details are available in Appendix C.
@@
Table 2: Comparison of classification performance using synthetic data generated by different LLMs with our method. We report the average performances of five runs of gradient boosting classifier. #syn denotes the number of synthetic samples added to the original dataset.
@@ 3.2 Machine learning classification performance using the synthetic data
We evaluate the quality of the synthetic data samples by assessing the machine learning classification performance when the synthetic data are added to the original dataset. Here, our method utilizes the GPT-3.5-turbo model. As shown in Table 1, our method achieves state-of-the-art F1 scores and balanced accuracy across all six datasets, surpassing the baselines that require model training. Notably, our method is the only one that consistently outperforms the original data in both F1 score and balanced accuracy across all six datasets. For a machine learning model to perform well in classification, the correlation between input data and labels in the training data must be precise. Thus, the consistent improvement across six datasets demonstrates that the data generated by our method aligns well with the class labels. Other methods often result in worse performance than the original data, indicating that the data generated by baselines disrupts the original data distribution. These findings underscore the robustness and effectiveness of our method, affirming its superiority over baselines, even for challenging datasets.
Moreover, baselines such as GReaT and TabDDPM exhibit significantly lower sensitivity compared to specificity, particularly for Travel and Income. Their high balanced accuracy is due to high specificity, but their severely low sensitivity indicates a failure to generate appropriate minority class data. In contrast, our method significantly improves sensitivity and achieves a balance between sensitivity and specificity while also attaining the best-balanced accuracy and F1 score compared to the baselines. For example, in Travel, our method achieves a sensitivity of 78%, which is 19%p higher than the second-best method, GReaT, at 58.8%. In Income, our method shows 66.45% sensitivity, surpassing the second-best score of 60.69%. These results highlight the effectiveness of our method in accurately representing minority classes in tabular datasets.
Overall, our findings demonstrate that our approach is generally applicable across various imbalanced tabular datasets and excels in generating high-quality samples that improve ML classification performance. Further analyses exploring the impact of augmenting the original dataset for the minority class (Table 6) and replacing the original dataset with synthetic data (Table 7) are available in Appendix A.
@@ 3.3 Open-source LLMs
Table 4: Comparison of token usage and generation efficiency across ablated methods on the Income dataset. Results are based on 100 inferences, each with 20 random input samples. Input tokens indicates the number of tokens required in the LLM prompt for a fixed number of input samples. Output samples shows the average number of synthetic samples generated per iteration. Success rate measures the ratio of inferences that generate at least one valid data sample.
@@
Table 5: Comparison of classification performance on the Sick dataset for task specification elements in prompt design. Results are averaged across XGBoost, CatBoost, LightGBM, and gradient boosting classifier. Methods marked with an asterisk (*) use the prompt designs proposed in the respective papers. #syn denotes the number of synthetic samples added to the original dataset.
@@ 3.4 Exploring optimal prompt design through ablation studies
We investigate key prompt design choices and validate our method through extensive ablation studies across multiple datasets from diverse domains, focusing on ML performance and generation efficiency, as detailed in Tables 3 and 4, respectively. Given a fixed number of input samples, we evaluated (1) the number of input tokens required, (2) the number of valid generated samples, and (3) the generation success rate. Our results indicate that CSV-style prompting generally outperforms sentence-style prompting in ML performance with the same number of input samples. While generating data for one class at a time yields good results for the Sick and Income datasets, it results in a low F1 score of 58.67% on the Travel dataset. However, generating data for multiple classes simultaneously with balanced samples improves this score to 70.37%. Furthermore, applying unique variable mapping consistently enhances F1 scores on the Sick dataset. The combined approach of random mapping and grouping achieves higher F1 score of 91.95% compared to 86.60% with random mapping but without grouping, underscoring the positive impact of this combined approach. These findings indicate that prompt design significantly influences the quality of generated data, emphasizing the need for carefully tailored prompts to generate high-quality synthetic tabular data.
@@
For the Travel and Income datasets, performance without grouping is slightly better than with grouping, but the differences are within 0.5%p, making their performance on par. However, these show significant differences in LLM generation efficiency. Without grouping, an average of 0.88 samples are generated per iteration, whereas grouping increases this to an average of 8.6 samples per iteration. Repeating the set twice within the prompt further improves generation efficiency, boosting the average output to over 10 samples and the success rate to 94%, by creating input patterns that LLMs can mimic to generate data. Unique variable mapping also significantly enhances LLM generation efficiency and success rates, as detailed in Table 8 of Appendix A.3.
Figure 4: Difference between Cramér’s V correlation matrices of real and synthetic datasets for categorical variables in the Travel dataset. More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue. Black indicates where the correlation is not measured since only one unique value is generated for those variables.
@@
Furthermore, we conduct an ablation study on the task specification elements, as shown in Table 5. Our approach of triggering completion significantly enhances overall ML performance compared to the instruction-based prompting methods. Interestingly, adding class distinctions within the prompt achieves the best sensitivity, indicating that the distinction facilitates better generation for minority classes. Providing variable descriptions improves balanced accuracy but reduces the F1 score, suggesting that this element may be optional rather than a definitive improvement. A comparison of generation efficiency using these methods is provided in Table 9 of Appendix A.3.
3.5 Analysis of feature correlation
@@
We compare the feature correlations of the categorical variables between original and synthetic data, as illustrated in Fig. 4. Our synthetic data generated with Mistral and Llama2 exhibit the closest feature correlation with the original data for both minority and majority classes, outperforming the baselines and ablated versions. Additionally, our grouping method significantly contributes to preserving feature correlation for the minority class. These findings indicate that our method generates data that more accurately matches each class, particularly for underrepresented classes, compared to other methods. Additional results on the Sick dataset are available in Fig. 5 of Appendix A.3.
4 Limitations and future work
When the training dataset is large and cannot be fully included in the LLM prompt due to token size limitations, only a subset of the data can be used as examples for generating samples. If these prompt samples do not fully represent the original data distribution, the generated data may be incomplete and of low quality, as LLMs are limited to producing data based only on the patterns present in the input. To overcome this, our method employs multiple rounds of random sampling with replacement to create a synthetic dataset that more comprehensively represents the original distribution, resulting in improved machine learning classification performance (a comprehensive analysis is provided in Appendix B.1). However, this approach still carries the risk that samples may not fully capture the original data distribution. Future research will focus on developing methods to identify key examples that more accurately represent the entire dataset.
@@ 5 Conclusion
This study demonstrates the effectiveness of using LLMs for synthetic tabular data generation. We introduce EPIC, a simple yet effective solution for generating realistic, high-quality data without additional training, specifically designed to address class imbalance. Our method demonstrates significant improvements over state-of-the-art generation models across six real-world public datasets, generating data with highly accurate feature correlations and significantly improving ML classification accuracy for minority classes. These results underscore the substantial impact of our approach in real-world applications, making a significant contribution to the field of tabular data research.
Acknowledgements
This work was supported by Institute for Information & communications Technology Promotion(IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program(KAIST)), Electronics and Telecommunications Research Institute(ETRI) grant funded by the Korean government (24ZB1220, Fundamental Technology Research for Human-Centric Autonomous Intelligent Systems), and the National Supercomputing Center with supercomputing resources including technical support (KSC-2024-CRE-0028).
@@ Acknowledgements
  • Appendix A provides additional experimental results on the machine learning classification for imbalanced data, ablation study, and feature correlation analysis.
  • Appendix B studies the impact of sample size on the machine learning performance and comprehensively analyzes the experiments conducted on the toy dataset.
  • Appendix C provides comprehensive information on the datasets, baselines, implementation details, and prompt design and examples.
  • Appendix D provides related work.
  • Appendix E discusses the broader impacts of our research.
  • Appendix F presents confusion matrix results and the complete results of Table 1.
@@ Appendix A Additional experimental results
Our experiments with synthetic tabular data for imbalanced classes follow three main approaches: (1) Adding synthetic data to both minority and majority classes of the existing dataset and evaluating machine learning classification performance. The results of this approach are presented in Table 1 of the main manuscript. (2) Adding synthetic data only to the minority class, with detailed results provided in Appendix A.1. (3) Evaluating classification performance using only synthetic data without the original dataset, as discussed in Appendix A.2. Moreover, Appendix A.3 presents further ablation study results, and Appendix A.4 provides a feature correlation analysis for the sick dataset.
Table 6: Comparison of binary classification performance when augmenting the minority class with synthetic data to balance class sizes. Average performance of the gradient boosting classifier is reported over five runs. #syn denotes the number of synthetic samples added to the original dataset.
@@
A.1 Comparative analysis of augmenting the original dataset for the minority class
One of the major advantages of our method is its ability to generate high-quality synthetic data for minority classes, even with limited samples. Similar to oversampling techniques like SMOTE, we add synthetic dataset only to the minority class and evaluate machine learning classification performance. To balance the ratio of majority to minority classes, we add only minority class samples to the original data and then measure the classification performance of the machine learning model. For this task, we additionally investigate the effectiveness of our method compared to SMOTE and SMOTENC [5]. As shown in Table 6, adding minority class samples generated from our method leads to higher F1 scores and balanced accuracy on both Travel and Income datasets, compared to when samples generated by other baselines are added. In the Travel dataset, where the number of minority class samples is substantially low, baselines, especially CopulaGAN and CTGAN, demonstrate markedly low sensitivity, indicating their failure to accurately learn the minority class data distribution. In contrast, our method demonstrates its efficacy in producing high-quality data for minority classes on the Travel dataset. Compared to the Original, it significantly improves the F1 score by 12.13%p, underscoring its superior performance of our approach.
A.2 Comparative analysis of replacing the original dataset with synthetic data
Table 7: Comparison of classification performance using only synthetic data for training. Average performance of the gradient boosting classifier is reported over five runs.